OpenStudy (anonymous):

A circle has its center at (-2, -3) and a radius of 3 units. What is the equation of the circle? what formula would i have to use to solve this. or better yet how would i solve this. Thanks.

OpenStudy (anonymous):

\[ f(x,y)=(x-u)^2+(y-v)^2 = r^2 \] where (u/v) are the xy-respective coordinates of the center and r is the radius.

OpenStudy (asnaseer):

you might also find this site useful: http://www.mathsisfun.com/algebra/circle-equations.html

OpenStudy (anonymous):

Thanks.

OpenStudy (asnaseer):

yw :)